OpenStudy (anonymous):

what is the int of tan2x

OpenStudy (anonymous):

write as \[\int\frac{\sin(2x)}{\cos(2x)}dx\] then a u- substitution will give it to you

OpenStudy (ash2326):

integral of tanx = log (sec x) so tan 2x ---> 1/2log (sec 2x)